OpenStudy (anonymous):

Jerome left town driving north at 60 mph. Anne left 2 hours later driving the same route at 76 mph. How many hours after Jerome left will it take for Anne to catch up to him? Round to the nearest tenth if necessary. Answer 7.5 hours 5.6 hours 9.5 hours 8.1 hours

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

@MathLegend help me out!

OpenStudy (anonymous):

Someone?

OpenStudy (anonymous):

About 7.5

OpenStudy (mathlegend):

@ktnguyen1 explanation please. :)

OpenStudy (anonymous):

d: distant d/60-d/76=2 Solve for d=570 t=570/76=7.5